### v3

# Michael Ong Dept. of Linguistics

Abstract
We introduce the Faetar Automatic Speech Recognition Benchmark, a benchmark corpus designed to push the limits of current approaches to low-resource speech recognition. Faetar, a Franco-Provençal variety spoken primarily in Italy, has no standard orthography, has virtually no existing textual or speech resources other than what is included in the benchmark, and is quite different from other forms of Franco-Provençal. The corpus comes from field recordings, most of which are noisy, for which only 5 hrs have matching transcriptions, and for which forced alignment is of variable quality. The corpus contains an additional 20 hrs of unlabelled speech. We report baseline results from state-of-the-art multilingual speech foundation models with a best phone error rate of 30.5%, using a pipeline that continues pre-training on the foundation model using the unlabelled set.
Index Terms:
Speech Recognition, Low Resource, Faetar, Self-Supervised, Unsupervised
I Introduction
Speech technology for low-resource languages has been accelerated by foundation models trained multilingually or on related languages. However, for the most part, demonstrations of the benefits of pre-training or cross-language transfer approaches for low-resource automatic speech recognition (ASR) have been broad-coverage evaluations aggregated over many languages and corpora at once [1]. This masks vast performance differences. The value of human linguistic diversity makes it important for the next era of speech recognition to deal with limited data, and speech in difficult conditions.
We introduce a new benchmark corpus to the community in the form of a challenge. The corpus is of Faetar, a variety of the Franco-Provençal language which developed in isolation in Italy, far from other speakers of Franco-Provençal, and in close contact with Italian and its dialects [2, 3]. Faetar has less than 1000 speakers around the world, in Italy and in the diaspora [4]. The benchmark data represents the majority of all archived speech recordings of Faetar in existence, and it is not available from any other source.
The language has no standard written form. However, as it is endangered, preservation, learning, and documentation are a priority for many community members. As such, automatic speech recognition would provide the valuable ability to transcribe and index Faetar recordings. We believe this situation is typical of many languages, as is the quantity and quality of the data set, which consists of around five hours of labelled speech (transcribed quasi-phonetically), and around 20 hours of unlabelled speech. All come from field recordings, generally noisy, of semi-spontaneous speech.
We believe exploration and advances can be made quickly by drawing the community’s attention to a common problem. As data of this kind is typical in many languages (noisy recordings, with only a few hours transcribed), we believe important, generalizable lessons can be drawn from comparing approaches on this data set. The benchmark, including baseline and evaluation code, is freely available subject to an online data-sharing agreement. [^1] https://perceptimatic.github.io/faetarspeech We release the benchmark in two phases: an embargoed phase, during which the reference transcriptions for the test set are only available to the organizers. This period will last roughly eight months and is intended to allow us to generate interest in the task by proposing it in the form of a challenge. Following this, the entire benchmark will be released.
II Related Work
Low-resource ASR benchmark tasks began with IARPA Babel, [^2] https://www.iarpa.gov/index.php/research-programs/babel, last accessed April 8, 2024. which aimed to facilitate the rapid development of robust ASR and keyword-spotting systems on languages with limited transcribed data [5]—around 200 hours per language. In 2020 and 2021, Babel data were repurposed for the NIST OpenASR challenges, with similar goals [6, 5]. In 2018, a low-resource speech recognition challenge for Indian languages was held [7], using about 40 hours of labelled data per language, and, in 2022, an ASR challenge for Hindi was proposed, using about 100 hours of labelled data [8].
While all of the languages in the above challenges are severely more resource-limited than English, substantial additional speech and textual data can be found for all of the languages used in the above challenges [9, 10, 11, 12, 13, 14]. Large collections of unlabelled data can boost performance substantially [15]. Thus, in this sense, the Faetar corpus is very different from corpora used in previous challenges, as there is almost no additional publicly available data, either speech or text, that challenge participants could make use of. Furthermore, the amount of labelled data is substantially smaller than most challenges. Previous challenges also generally use higher-quality recordings. While other endangered languages such as Yoloxóchitl Mixtec and Sierra Totonac are similarly under-resourced [16, 17, 18], these languages have not been explicitly presented to the community as benchmark tasks.
Recent benchmark evaluations have been developed for multilingual and low-resource speech recognition, focusing on breadth over depth, evaluating on as many languages as possible [19, 14, 20, 21, 1]. Evaluating on a wide breadth of languages represents a great advance over the field’s traditional attention to a few languages. However, aggregate scores mask radical variation across languages and corpora in how well state-of-the-art approaches work, and more generally which strategies work best [1, 21, 15, 22]. The Faetar benchmark complements breadth-first benchmarks with the depth of a single language, challenging researchers on a single, difficult ASR task to compare approaches in a controlled fashion.
III Materials
III-A Language
Faetar (pronounced [fajdar]) is spoken in Faeto, in the Apulia region of southern Italy, and is very similar to the language spoken in the neighbouring town of Celle St. Vito. The population of the two towns is less than 1000, and Faetar is threatened by population loss as well as language shift to Italian. Because of contact with Italian, the language has absorbed many Italian loan words, and, because of bilingualism, Faetar speakers often code-switch, using Italian words in Faetar. The language is very different from Italian both in its structure and in its base vocabulary, as it is a variety of Franco-Provençal, a member of the southern group of Gallo-Romance languages (French is in the northern group, which are quite different in pronunciation and vocabulary). Franco-Provençal was brought to Apulia by a group of speakers some time between the 12th and 15th centuries (according to tradition, soldiers fighting for Charles of Anjou), and the community has since been entirely isolated from the rest of the Franco-Provençal speaking community [3, 4]. There is also a Faetar-speaking diaspora, including in the Greater Toronto Area in Canada [23], where some of the recordings were made.
[…]
{talltblr}
[theme=nocaption] colspec=lcccccccccccccccc, cell2-Z2-Z=, vlineeven[2-Z], vline1, cell1even[2-Y]=c=2c, hlines, column2-Z=wd=2em,c, columns=colsep=1pt, vlineodd[3-Z]=dashed,gray6, & Bilabial Labiodental Alveolar Postalveolar Palatal Velar
[…]
TABLE: The phonemic inventory of Faetar (minus the phonemic geminate consonants). See the text for discussion of additional phones used in the transcriptions.
III-B Data
Data were extracted from the Faetar collection of the Heritage Language Variation and Change in Toronto (HLVC) corpus [24]. The corpus contains 184 recordings of native Faetar speakers collected in Faeto between 1992 and 1994 (the Homeland subset) and 37 recordings of first- and second-generation heritage Faetar speakers collected in Toronto between 2009 and 2010 (the Heritage subset) [24]. Some of the recordings were interviews (Interview), designed to elicit spontaneous speech, and, in others, participants were asked to describe scenes and objects from a picture book (Words). The recordings were saved at a 44.1 kHz sampling rate (in the case of the Homeland recordings, this is the sampling rate of the digitization from analog cassette tapes). Recordings can be noisy, with considerable background noise, back-channels, and interruptions.
The source data set consisted of long-form recordings, of which 68 of the Homeland recordings, and 26 of the Heritage recordings, had been at least partially phonetically transcribed at the utterance level. Some of the transcriptions (mostly those in the Heritage subset) were extracted from ELAN annotation files, which had utterance-level alignments [25], while the remainder were extracted from Microsoft Word files, by manually mapping numerical values from each of the two fonts used (PalPhon or IPAPhon) to modern UTF-8 encodings, and removing anything that did not look like a transcription (English glosses, identifiers, etc.).
For the portion which was transcribed but not aligned, we adapted the JHU Arabic MGB-3 recipe for segmentation and alignment [26], built on Kaldi [27]. This splits recordings and recording-level transcriptions by i-vector speaker diarization, splits each diarized segment into smaller, overlapping segments, aligns each overlapping segment to the text (allowing for minor changes to the text) with a seed ASR system, and stitches the overlapping segments back together using string matching. We trained the seed ASR system on the aligned part of the data. To generate our automatic alignments, we used a monophone, speaker-independent Gaussian mixture model-Hidden Markov model (GMM-HMM) system. We found this worked better than triphone or speaker-dependent models, presumably because of the very limited aligned data we had available for training. We then used PyAnnote 3.0 [28, 29] to adjust utterance boundaries according to voice-activity detection and label them with speaker diarization. Because the pipeline is error-prone, we manually threw out utterances whose boundaries were clearly misaligned or whose transcriptions were very clearly wrong. On the test set (see below), we made a second, more rigorous, manual pass to correct alignments.
After obtaining time-alignments for the whole data set, further filtering was performed to remove the interviewer’s speech (generally clearly marked), to remove utterances with duration less than 500 ms, and to remove utterances in Italian or English. While code-switching is common in Faetar speech (and thus, a recognizer should deal with it) removing utterances with substantial parts of Italian or English did not result in a major loss of data. As such, we take it that the (simpler) problem we propose, which consists of transcribing Faetar with minimal code-switching, is a reasonable approximation to the real problem. For the purposes of ASR research, it will be easier to interpret the results in a more homogeneous corpus. Nevertheless, as Faetar has been in contact with Italian for a long time, and substantially longer than it has been in contact with English (English only appears in the Heritage portion of the corpus), we took a more lenient approach to identifying Italian utterances. To identify English utterances, the presence of a single English word, drawn from a closed list that were clearly not words in Faetar, was sufficient to mark an utterance as English. On the other hand, for Italian, we required that there be three words in a row, again from a closed list.
We split the aligned data into train, dev, and test, ensuring a reasonable balance between male and female speakers, and between Homeland and Heritage subsets, in both train and test. We do not include the Words subset in the dev or test splits, as we take it that this is too easy an evaluation: many of the utterances consist of individual words, often the same word. Anticipating compatibility with the ML-SUPERB benchmark [1], we also construct 1h and 10min subsets of train. We distribute the remainder of the data, for which we did not have transcriptions, or for which the (long-form) file could not be time-aligned, as an unlabelled set, after obtaining VAD and speaker diarization using PyAnnote 3.0. Because we determined that many of the shorter utterances extracted in this way were not speech, we discarded all unlabelled utterances less than 1.5 seconds. Table I shows the distribution of data in the corpus.
TABLE I: Distribution of data in (hh:)mm:ss.
Homeland Heritage Total
M F M F
dev Int. 8:43 0:31 0:00 2:35 11:49
test Int. 13:10 13:33 10:54 9:16 46:54
train Words 42:31 58:29 51:44 9:07 4:30:17
Int. 15:58 27:56 32:49 14:41
1h Int. 11:08 13:44 32:50 0:51 58:34
10min Int. 0:00 4:25 5:00 0:24 9:49
unlab Words 7:29:11 0:05:24 19:55:21
Int. 8:23:18 0:51:31
Mixed 2:12:47 0:53:09
IV Evaluation
We take the most relevant measure of performance to be the phone error rate (PER) on the test set, calculated by aligning reference and hypothesis transcripts by the unit-cost Levenshtein algorithm. Given all utterances $u$ in a partition $\mathcal{U}$, the number of phones in a reference transcript $n_{u}$, and the number of errors in a model’s hypothesis transcript $e_{u}$, the PER on $\mathcal{U}$ is defined as:
$\text{PER}_{\mathcal{U}}=\frac{\sum_{u\in\mathcal{U}}e_{u}}{\sum_{u\in\mathcal{U}}n_{u}}\times 100\%.$ (1)
The phonetic transcriptions contain spaces between words, making both the word error rate (WER) and character error rate (CER) possible alternative evaluation metrics. Both are also calculated by our evaluation tool, but WER is much too severe as the transcriptions include substantial phonetic variation that would not ordinarily be present in orthographic transcripts (see above). PER and CER (which differ by the exclusion or inclusion of spaces, and by the treatment of affricates and long phones as a single phone or two characters) do not differ by more than a few percentage points in our observations. We chose PER (no spaces, affricates and long phones treated as single units) to avoid making the task unnecessarily difficult.
V Baseline Experiments
We consider unconstrained and constrained baseline systems, where by “constrained” we mean that the model is not trained on data sets other than train (thus precluding pre-trained models trained on other languages, as well as unlab). Considering constrained approaches separately allows us to focus on questions of model architecture, rather than on pre-training. As far as constrained models go, as discussed in Section III-B, force-alignment of transcriptions to audio was performed with Kaldi [27], initially using a monophone, speaker-independent HMM-GMM model, and then refined at the end using a triphone, speaker-dependent model; both were composed with a 5-gram, character-level, modified Kneser-Ney language model [30] trained on the (post-alignment) transcriptions from the train partition. We report the performance for these models in Table II, $\pm$ the half width of a 95% bootstrap confidence interval calculated using $K=10000$ samples.
TABLE II: Phone error rates (PERs, %) for each baseline (rows) and test partition (columns), $\pm$ the half width of a 95% bootstrap confidence interval.
Train Constr. Unlab. Train Dev Test
  HMM-GMM Mono +5gr Train 47.1 ±$\pm 0.6$ 65.9 ±$\pm 1.8$ 62.6 ±$\pm 0.8$
HMM-GMM Tri +5gr Train 9.6 ±$\pm 0.4$ 58.2 ±$\pm 2.0$ 56.7 ±$\pm 0.9$
ESPnet Train 36.0 ±$\pm 0.5$ 43.0 ±$\pm 2.8$ 35.9 ±$\pm 0.8$
  MMS Fine-tuning (FT) Train 34.4 ±$\pm 0.5$ 39.9 ±$\pm 2.8$ 33.0 ±$\pm 0.8$
mHuBERT-147 FT Train 26.6 ±$\pm 0.5$ 41.1 ±$\pm 2.6$ 33.6 ±$\pm 0.8$
  MMS Pre-training (PT) + FT Train + Unlab 31.4 ±$\pm 0.5$ 38.7 ±$\pm 2.7$ 31.5 ±$\pm 0.8$
MMS Self-training (ST) Train + Unlab 31.4 ±$\pm 0.5$ 36.6 ±$\pm 2.7$ 31.0 ±$\pm 0.8$
MMS PT + ST Train + Unlab 30.0 ±$\pm 0.5$ 36.4 ±$\pm 2.7$ 30.5 ±$\pm 0.8$
  ESPnet 1hr 39.3 ±$\pm 0.5$ 44.4 ±$\pm 2.8$ 37.4 ±$\pm 0.8$
ESPnet 10min 48.2 ±$\pm 0.5$ 50.2 ±$\pm 2.4$ 45.1 ±$\pm 0.8$
Alternatively, we train using the ML-SUPERB recipe [1] from ESPNet [31] using filterbanks, keeping all hyperparameters as-is except to set the effective batch size to 4. The model consists of a single convolutional layer followed by two transformer layers, trained using CTC. This was substantially better than the HMM-GMM model. [^3] We also trained an LSTM model with CTC. Its performance was not substantially better than the HMM-GMM models, so we do not report it here. Decoding with a 5-gram language model trained on train did not improve PERs, presumably because the neural model already exhausted what can be learned from train. For comparison with ML-SUPERB, we also include results for the restricted 1hr and 10min training conditions.
In the unconstrained condition, we fine-tune existing multilingual models MMS [32] and mHuBERT-147 [33] on the train set. Both are masked prediction models; MMS is a wav2vec 2.0 model that takes a language-specific adapter, head, and fine-tuning strategy. Rather than initializing the adapter layers from scratch, we initialize the adapters and head with a model fine-tuned on Italian. mHuBERT-147 takes an iterative refinement approach, training first on classification according to classes derived from k-means clustering on spectral features, then, after the first iteration, on targets derived from clustering the internal representations of the previous one. As [33] obtained best results with three iterations, we fine-tune the third iteration. For both models, we use a linear schedule with peak learning rate of $1e^{-5}$ for 200 epochs of which ten are warmup, dropout of 10%, and effective batch size of 8. Both give substantially better results than ESPnet, and the differences in performance between the two are minor, with MMS having a slight advantage (within the overlap of the confidence intervals).
The data includes an unlabelled subset. We show two possible approaches to using it. First, continued pre-training of a multilingual model using Faetar audio for self-supervised training prior to fine-tuning. We use the entire unlab set and continue pre-training of MMS (effective batch size 32; 300000 steps $\approx$258 epochs, of which the first 30% are warmup in a linear scheduler; peak learning rate $3\times 10^{-4}$), prior to fine-tuning as before. As can be seen from Table II, this leads to a large performance boost. Alternatively, we attempt self-training, decoding unlab using the fine-tuned (base) MMS model, then adding the resulting decoded speech to a second round of fine-tuning. Finally, we combine both approaches by fine-tuning the model resulting from continued pre-training instead of the base model. Table II demonstrates that both of these approaches improve performance, and that the results are cumulative, leading to a best PER of 30.5% when combining both approaches (although the pairwise differences are within the overlap of the confidence intervals again). In sum, the unlabelled data set shows potential for improving ASR performance.
VI Limitations
Many languages are in a similar situation to Faetar: very close to no speech or textual resources, and what speech resources there are are noisy and only partially transcribed. However, each corpus is different (e.g., for many languages, far less recorded speech exists than this), and there is no guarantee that the lessons learned from this benchmark will generalize to other settings.
The fact that Faetar has no standard orthography does not detract from the utility of ASR: a major application of ASR for minoritized and endangered languages is in indexing audio recordings and in generating textual resources from them (that may, for example, later be used to build educational materials), in addition to linguistic research. Document collections can be built without a standard orthography. Nevertheless, to be useful, the language community eventually needs to be able to use the transcription system.
Lessons learned through this benchmark are also relevant to speech-only applications (speech-to-speech translation, spoken language generation and dialogue), but with this limited amount of data, the feasibility of these traditionally more resource-intensive tasks remains to be seen.
VII Summary of Contributions
We have presented the Faetar Automatic Speech Recognition Benchmark, an open benchmark corpus with evaluation and baseline code aimed at very under-resourced languages. We demonstrate that out-of-the-box application of state-of-the-art multilingual foundation models yields an improvement over spectral features, and that using unlabelled data yields further improvements. The fact that we obtain improvements on the constrained ASR task by varying models suggests that there may be further improvements to be made to the model architectures as well. Nevertheless, in all cases, the baseline PERs would generally be unacceptable without manual correction. We leave additional approaches (e.g., improved speaker modelling, speech enhancement, hybrid models) for the community to explore, and release the benchmark with the aim of encouraging careful attention to corpora and languages like Faetar, deepening the field’s commitment to the promise of universal speech technology.